# Image Classification in High-Energy Physics: A Comprehensive Survey of Applications to Jet Analysis

### v2

@@ 5.4 Adversarial training-based methods
GANs in image processing enhance creativity and realism by generating new images through a dynamic interplay. The generator creates images, while the discriminator evaluates and refines them, enabling tasks like image-to-image translation, style transfer, and data augmentation with unparalleled versatility [60, 53]. GANs are powerful tools for Jet image classification in particle physics. They create realistic Jet images, enabling robust testing of classification algorithms. GANs enhance the accuracy of identifying particles and contribute to breakthroughs in HEP research. However, the authors in [61] employed another technique for adversarial training for physics object identification and decreased the effect of simulation-specific artifacts. They systematically distorted inputs that have been generated with fast gradient sign method (FGSM) adversarial attack technique, this latter altering model predictions using gradient information. The method showed how model performance and robustness are related. They explored the trade-off between performance on unperturbed and on distorted test samples, investigating ROC curves and AUC scores for the used discriminators. Similarly, in [62], the paper investigates the loss manifold of a Jet tagging algorithm concerning input features on nominal and adversarial samples. Discrepancies in flatness reveal differences in robustness and generalization. The study suggests refined training approaches through macro-scale loss manifold exploration for two features and devising attacks that maintain the gradient’s directionality. This leverages acquired insights for enhanced object identification in particle physics.
@@ 5.5 RNN-based methods
Various types of RNNs such as bidirectional RNNs (BRNNs), LSTM, and gated recurrent units (GRUs) differ in architecture at the cell level within the RNN layer. BRNNs propagate information in both forward and backward directions, influencing predictions by surrounding words. LSTM tackles vanishing gradients with inner cells containing input, output, and forget gates, regulating information flow. GRUs-based networks address short-term memory issues with reset and update gates controlling information utilization akin to LSTM gates [63, 52]. Recursive neural networks (RecNNs), are designed to operate on hierarchical or tree-structured data, where the relationships between elements are defined by a recursive structure. Instead of processing sequences sequentially with temporal dependencies, like RNNs, RecNNs recursively apply the same neural network operation to combine representations of child nodes to produce a representation of their parent node, traversing the hierarchical structure. In light of this, the authors in [64] investigate RecNNs for quark/gluon discrimination. Results indicate RecNNs outperform baseline, boosted decision tree, in gluon rejection rate by a few percent. Even with minimal input features such as $p_{T},\eta,\phi$, RecNNs yield promising results, suggesting tree structure contains essential discrimination information. Additionally, rough up or down quark jet discrimination is explored. In [63], a neural network was created specifically for Jet binary classifying . The network comprises two hidden layers employing recurrent cells, with a structure consisting of 25 LSTM cells and utilizing a tanh activation function at its core.
5.6 MLP-based methods
MLP is an artificial neural network composed of multiple layers of nodes, including an input layer, one or more hidden layers, and an output layer. Each node in one layer is connected to every node in the subsequent layer. MLP can handle complex nonlinear relationships between input and output data, making them suitable for various tasks. MLPs are versatile, scalable, and can be trained using back-propagation, enabling them to learn from large datasets effectively and generalize well to unseen data. Kinematic parameters describe the motion of particles, including velocity, momentum ($p_{T,J}$) and trimmed jet momentum ($p_{T,J,trim}$), energy, jet mass $m_{J}$ and jet mass trimmed $m_{J,trim}$ , and angles of emission, commonly used in physics and engineering analyses. Chakraborty et al. in [65] employed both kinematics and spectral function, which typically refers to a function that describes the distribution of energy or momentum states of particles in a particular physical system, to feed MLP classifier as described in Figure 16. The authors aim is to trim/discard Jet that are unlikely to have originated from the process of interest (effects of background noise). This selective removal helps to improve the accuracy of measurements and analyses by focusing only on the most relevant particles within a jet.
Table 6: Summary of the performance of certain DL frameworks proposed for HEP. Only the best performance is reported in the case of multiple tests.
@@ 5.6 MLP-based methods
[34] SVM Simulated BIP features invariant under boosts for improved Jet tagging Acc= 92.7 Performance could be enhanced through comprehensive hyperparameter tuning. Yes [^12] https://zenodo.org/records/7271316
@@
[69] ANN Simulated Identifying boosted top quarks using pattern recognition through an ANN in HEP experiments. Eff= 60 It has 4% mis-tag rate. It exclusively utilizes hadronic calorimete (HCAL) data, though additional data, like sub-jet b-tags, are crucial for top tagging. No
[70] DNN Real data Enhancing jet reconstruction at CMS through deep learning FPR= 65 The computational costs, wnen employing the proposed model, have not been verified. No
@@
Abbreviations: Deep learning model (DLM); Best performance (BP); Project link availability (PLA); auto-encoder (AE).
6 Applications of AI-based Jet image
@@
Jet images processed through ML and DL techniques hold vast potential across various applications within the HEP domain, some of theme are already described in [17]. This section presents a comprehensive overview of cutting-edge work in this area, categorized into several key domains: Jet parameter scanning, event classification, Jet tagging, multi-Jet classification, energy estimation, and beyond [72]. The taxonomy of AI-based jet image applications is visualized in Figure 17, illustrating their scope and relationships. The section thoroughly reviews some applications conducted by researchers, while suggesting future directions for those not yet explored. Additionally, Table 6 provides a concise summary of performance metrics, limitations, online project availability, and results obtained across these applications, offering valuable insights into their efficacy and applicability.
Figure 17: Taxonomy of AI-based HEP applications using Jet images.
6.1 Parameters scan
@@
The utilization of a ML and DL models enables the comprehension and estimation of the correlation between the parameter space of the new physics model and the experimental physical observables. This facilitates the efficient constraint of the parameter space of the new physics model [17]. Given the sensitivity of the ATLAS experiment to exploring the parameters and event counts in a new physics, significant computing power is required to deduce the surviving regions of the parameter space of constrained minimal super-symmetric standard model (CMSSM) using Bayesian posterior probability and likelihood function ratio test. To mitigate computational demands, the study [73] utilizes a MLP as a regressor to learn the mapping from CMSSM model parameters $\theta$ to weak-scale super-symmetric particle masses $m$. The output of the SoftSusy physical package serves as the target output value of the neural network. Approximately 4000 sample points in the parameter space form the training set to train the regressor. With a given set of CMSSM parameters, this MLP model rapidly predicts the corresponding super-symmetric particle mass spectrum. This approach significantly accelerates the process compared to traditional methods. To identify the parameters of a new physics model, [74], a MLP was trained using 84 physical observables from the 14 TeV LHC as inputs, with the parameters of a super-symmetric model as the desired outputs. The study revealed that with a collider luminosity of 10 $fb^{-1}$, the CMSSM model’s parameters $M_{0}$ and $M_{1/2}$ could be reliably determined with just a 1% margin of error. With a collider luminosity of 500 $fb^{-1}$, additional model parameters such as tan $\beta$ and $A_{0}$ could also be accurately estimated. In contrast, the conventional approach of minimizing $\chi^{2}$ yielded comparatively inferior results. Generating collider event samples at the LHC through Monte Carlo simulation can be a time-intensive process. While a rapid detector simulation typically requires only a few minutes, a comprehensive simulation using the GEANT4 framework, as employed by experimental groups like ATLAS and CMS, may necessitate several days. To address this, parallel full detector simulations are conducted using four parameters , namely common scalar mass ($m_{0}$), the universal gaugino mass ($m_{1/2}$), the trilinear coupling term ($A_{0}$), and the the ratio of the vacuum expectation values ($\tan\beta$) [75]. Two ML models, namely the MLP and SVM, were employed to learn the correlation between the number of signal events and the CMSSM parameters. Results showed that the accuracy of predicting the likelihood function could reach several percent with just 2000 training samples.
@@ 6.2 Jet classification and tagging
Despite treating Jets as images in the calorimeter and exploiting the benefits of DNNs in image classification for improved Jet substructure detection, these approaches encounter hurdles. Challenges such as Jet image sparsity and potential precision loss arise from constructing Jet images through pixelation or creating advanced Jet features. In this study [42], a sequential method is employed, utilizing an ordered sequence of jet constituents as inputs for training. Unlike many prior methods, this approach avoids information loss during pixelization or high-level feature computation. The Jet classification technique achieves a considerable background rejection efficiency operating point for reconstructed jets with transverse momentum ranging from 600 to 2500 GeV. Moreover, it remains unaffected by multiple proton-proton interactions at levels anticipated during Run 2 of the LHC.
Particles generated in a collider with significant center-of-mass energy typically exhibit high velocity. As a result, their decay products tend to align closely, leading to overlapping jets. It is crucial in collider data analysis to discern whether a jet originates from a solitary light particle or from the decay of a heavier particle. Traditional approaches rely on manually crafted distribution features based on energy deposition in calorimeter cells. However, due to the intricate nature of the data, ML techniques have proven more efficient than human efforts for this task [76]. In [77], the Jet image concept treats the detector as a camera, capturing Jet energy distribution in calorimeters as a digital image. This enables Jet tagging as a pattern recognition task, utilizing ML, like Fisher classification, to differentiate between hadronic W boson decay and Jets from quarks or gluons. Monte Carlo simulation shows superior discrimination compared to traditional methods, offering insights into Jet structure. In [55], CNNs improve tagging by treating Jet energy distribution as an image, using channels for features like particle momentum and count. Results show CNNs can surpass traditional methods, providing reliable insights from collider simulation data despite variations in event generators. However, CNNs demonstrate a lack of sensitivity to quark/gluon Jets from different generators, akin to conventional Jet measurements. Moving on, in [78], Jet tagging is performed using RNN, leveraging the similarity between Jet clustering and natural language structure. Final-state particle four-momenta are treated as language words, and Jet clustering as grammatical analysis. RNN efficiently processes the tree-like Jet structures, enabling direct use of particle data regardless of count. This method yields higher data utilization efficiency and prediction accuracy than Jet image-based ML, extending to event classification. In [64], RNNs distinguish quark and gluon Jets, showing higher gluon suppression. Factors affecting RNN performance are explored, with preliminary quark tagging results. Numerous explorations for phenomena beyond the SM at the LHC depend on top tagging techniques that distinguish between boosted hadronic top quarks and the more prevalent jets that originate from light quarks and gluons. The HCAL essentially captures a "digital image" of each jet, where the pixel brightness represents the energy deposited in HCAL cells. Therefore, top tagging is essentially a matter of recognizing patterns. The work in [69] propose a novel top tagging algorithm based on an ANN, a popular pattern recognition approach. The ANN is developed using a substantial dataset of boosted tops along with light quark/gluon jets and is subsequently evaluated on separate datasets. In Monte Carlo simulations, particularly within the 1100-1200 GeV range, the ANN-based tagger demonstrates outstanding efficacy.
@@
Multi-jet classification is a a key task in particle physics aimed at distinguishing between events with varying numbers of jets. Using ML techniques, such as DNNs, researchers develop classification models to accurately identify these events. Achieving high classification accuracy is crucial for understanding fundamental particle interactions and discovering new physics phenomena in experiments like those conducted at the LHC. The work in [16] present an application of scalable DL to analyze simulation data from proton-proton collisions at 13 TeV in the LHC. The researchers developed a CNN model which utilizes detector responses as two-dimensional images reflecting the geometry of the CMS detector. The model discriminates between signal events of R-parity violating super-symmetry and background events with multiple jets resulting from inelastic QCD scattering (QCD multi-jets). With the CNN model, they achieved 1.85 times higher efficiency and 1.2 times higher expected significance compared to the traditional cut-based method. They demonstrated the scalability of the model at a large scale using high-performance computing (HPC) resources with up to 1024 nodes. The authors in [65] proposing an interpretable network for multi-jet classification using the jet spectrum, termed S2(R), derived from a Taylor series of an arbitrary jet MLP classifier function. The network’s intermediate feature is an infrared and collinear safe variables, named C-correlator, estimating the importance of S2(R) deposits at angular scales. It offers comparable performance to CNNs with simpler architecture and fewer inputs.

### v3

@@ 4.4 Adversarial training-based methods
[…]
@@ 4.5 RNN-based methods
[…]
Table 6: Summary of the performance of certain ML and DL frameworks proposed for HEP. Only the best performance is reported in the case of multiple tests.
@@
[…]
@@
[…]
[79] DNN Real data Enhancing Jet reconstruction at CMS through DL FPR= 65 The computational costs, wnen employing the proposed model, have not been verified. No
@@
Abbreviations: DL model (DLM); Best performance (BP); Project link availability (PLA); auto-encoder (AE).
4.6 GNN-based methods
[…]
GNN are neural networks designed for graph-structured data, learning node and edge representations while capturing complex relationships and dependencies within graphs for tasks such as classification and prediction. In the HEP context, the authors in [49] proposed the ParticleNet model (Figure 13). The architecture is a customized neural network that operates directly on particle clouds for Jet tagging. It uses dynamic graph CNNs to process the unordered set of constituent particles that make up a Jet. The architecture consists of three EdgeConv blocks, each with a different number of channels and nearest neighbors. EdgeConv block starts by representing a point cloud as a graph, whose vertices are the points themselves, and the edges are constructed as connections between each point to its K-nearest neighbors (KNN) points. The EdgeConv block then finds the KNN particles for each particle, using the "coordinates" input of the EdgeConv block to compute the distances. Inputs to the EdgeConv operation, the "edge features," are constructed from the "features" input using the indices of KNN particles. The EdgeConv procedure is executed using a three-layer MLP. Each layer is structured to include a linear transformation, succeeded by batch normalization, and subsequently a ReLU activation. Additionally, a shortcut connection is integrated into every block parallel to the EdgeConv operation, facilitating the direct passage of input features. An EdgeConv block is defined by two key hyper-parameters: the neighbor count $k$ and the channel count $C$, which respectively denote the number of neighbors to consider and the number of units within each layer of linear transformation. The EdgeConv blocks play a crucial role in learning the local features of the particle cloud and aggregating them into a global feature vector for the Jet. Following EdgeConv blocks, global average pooling aggregates particle features, leading to a 256-unit fully connected layer, ReLU activation, dropout, and a 2-unit softmax output for binary classification. The ParticleNet architecture achieves SOTA performance on two representative Jet tagging benchmarks and is improved significantly over existing methods.
Figure 13: The architecture of ParticleNet model suggested in [49].
Similarly, [50] proposed the equivariant graph neural networks (EGNN) model, which is a GNN architecture that is translation, rotation, and reflection equivariant (E(n)), and permutation equivariant with respect to an input set of points. It uses a set of filters that are equivariant to the action of the symmetry group, which are constructed using a combination of radial basis functions and Chebyshev polynomials. The EGNN algorithm possesses the same flexibility as the GNN technique, while also maintaining E(n) equivariance similar to the radial field algorithm. Additionally, it eliminates the requirement for computationally intensive procedures, such as spherical harmonics. The EGNN exceeds other equivariant and non-equivariant options while maintaining efficiency in terms of running time. Moreover, the EGNN approach demonstrates a 32% reduction in error compared to the SOTA method.
Figure 14: The architecture of EGNN model suggested in [50].
Another architecture called the LorentzNet is proposed in [31], which is based on the Lorentz group equivariant block (LGEB) block. The structure of LGEB consists of several layers, including Minkowski norm and inner product, sum pooling, a MLP, and a Clebsch-Gordan tensor product. The input of LGEB is a set of 4-momentum vectors, which are transformed by the Minkowski norm and inner product layer to obtain Lorentz-invariant geometric quantities. The sum pooling layer aggregates the geometric quantities to obtain a scalar representation of the input. The MLP layer is used to learn a nonlinear mapping from the scalar representation to a new feature space. Finally, the Clebsch-Gordan tensor product layer is used to combine the new feature space with the original input to obtain the output of LGEB. It is designed as a Lorentz group-equivariant mapping to preserve the symmetries of the Lorentz group, ensuring the model’s equivariance and universality.
The paper [53] introduced Clifford group equivariant neural networkss, a novel GNN framework designed to construct $O(n)$- and $E(n)$-equivariant models using Clifford algebra. CGENNs leveraged the geometric properties of Clifford algebras, such as the geometric product, to parameterize equivariant neural network layers. These layers operated on multivectors—structures encompassing scalars, vectors, and higher-dimensional geometric features—enabling symmetry-aware computations. Input point cloud included scalars (e.g., mass) and vectors (e.g., positions), embedded into multivector subspaces. CGENNs achieved SOTA performance across domains, including 3D $n$-body simulations and 4D Lorentz-equivariant tasks, and Jet tagging in HEP, outperforming models like LorentzNet and EGNN. However, their computational costs, due to complex geometric products, remained a challenge for scalability and real-time applications.
4.7 Transformer-based methods
Transformers are AI models using self-attention mechanisms to process sequential data, excelling in natural language processing [81], computer vision [82], and time-series tasks by capturing long-range dependencies and contextual relationships efficiently. Researchers in HEP have investigated transformers for the Jet tagging task. For example, [51] introduced a modified point cloud Transformer (PCT) for jet-tagging tasks in collider physics. The PCT leveraged self-attention layers and EdgeConv blocks to handle the unordered nature of particle data, ensuring permutation invariance. Jets were represented as point clouds with up to 100 particles, described by kinematic features such as momentum and particle types. The suggested PCT achieved SOTA performance, with an high AUC for both top tagging and quark-gluon classification, showing up to a 20% improvement in background rejection over models like ParticleNet. Despite its superior performance, the computational cost was significant, with 266M FLOPs, making real-time applications challenging.
In addition, the work in [52] proposed PartT, which is a new Transformer-based architecture for Jet tagging. Its main task is to identify the origin of a Jet of particles produced in HEP experiments. ParT makes use of two sets of inputs: (i) the particle input, which includes a list of features for every particle and forms an array, and (ii) the interaction input, which is a matrix of features for every pair of particles. ParT employs a novel pairwise multi-head attention (P-MHA) mechanism, which allows the model to attend to pairs of particles and learn their interactions. The P-MHA is more effective than standard plain multi-head attention. This assertion is substantiated when the pre-trained ParT models are fine-tuned on two widely adopted Jet tagging benchmarks, the quark-gluon tagging dataset and the binary classification dataset for identifying boosted $W$ bosons decaying to two quarks. The fine-tuning process involves training the ParT models on a smaller labeled dataset specific to each benchmark, which allows the models to learn the specific features and patterns relevant to each task. The fine-tuned ParT models achieve significantly higher tagging performance than the models trained from scratch and outperform the previous SOTA models, including ParticleNet and other Transformer-based models.
Moving on, [54] introduced the Lorentz geometric algebra Transformer (L-GATr), a versatile architecture designed for high-energy physics. L-GATr combined Lorentz-equivariant geometric algebra with attention mechanisms, enabling robust handling of particle physics data in four-dimensional spacetime. The architecture accommodated variable-length inputs, exploited Lorentz symmetry, and extended to generative modeling via continuous normalizing flows trained with Riemannian flow matching. It used Transformer-based layers with Lorentz-equivariant attention and normalization tailored to Minkowski space, processing particle data parameterized by type and four-momentum vectors. The evaluation employed metrics such as accuracy, AUC, background rejection rates, MSE, likelihood, and two-sample tests. L-GATr demonstrated competitive or superior performance compared to Lorentz-equivariant graph networks. However, it had computational overhead relative to standard transformers and left its potential for pretraining in HEPs unexplored. Similarly, more-interaction particle Transformer (MIParT) scheme [55] introduced the more-interaction Attention (MIA) mechanism to enhance Jet tagging by embedding detailed particle interactions. Based on the Transformer architecture, MIParT-L doubled the dimensions of interaction embeddings for large datasets while reducing model complexity, with 30% fewer parameters and 53% lower computational demands than its predecessor, ParT. Tested on top tagging and quark-gluon datasets, MIParT-L achieved nearly identical accuracy and AUC to leading models while improving background rejection by 25% and 3%, respectively. Fine-tuning on large pre-trained datasets further improved performance by 39% and 6%. Despite its efficiency, the interpretability of MIParT-L remained a challenge, limiting insights into its decision-making process. This trade-off underscored the computational of model efficiency and robust performance across diverse Jet tagging tasks.
5 Applications of AI-based Jet classification
@@
Jet images and PC processed through ML and DL techniques hold vast potential across various applications within the HEP domain, some of theme are already described in [18]. This section presents a comprehensive overview of cutting-edge work in this area, categorized into several key domains: Jet parameter scanning, event classification, Jet tagging, multi-Jet classification, energy estimation, and beyond [83]. The taxonomy of AI-based Jet image and PC applications is visualized in Figure 17, illustrating their scope and relationships. The section thoroughly reviews some applications conducted by researchers, while suggesting future directions for those not yet explored. Additionally, Table 6 provides a concise summary of performance metrics, limitations, online project availability, and results obtained across these applications, offering valuable insights into their efficacy and applicability.
Figure 17: Taxonomy of AI-based HEP applications using Jet images or PC.
5.1 Jet parameters scan
A parameter scan in HEP involves systematically exploring a wide range of values for the theoretical parameters that define a given model. These parameters often characterize the masses of new particles, coupling strengths, or other fundamental quantities hypothesized in extensions of the SM. By examining different combinations of these parameters, researchers aim to identify which sets are compatible with current experimental data or make predictions that can be tested in future experiments. This process helps narrow down the vast theoretical landscape to more plausible scenarios, guiding ongoing investigations and informing the design of new searches [84].
The utilization of ML and DL models enables the comprehension and estimation of the correlation between the parameter space of new physics models and the experimental physical observables, including signatures characterized by Jets, leptons, and missing transverse energy. This facilitates the efficient constraint of the parameter space of the new physics model [18]. Given the sensitivity of the ATLAS experiment to exploring parameters, event counts, and Jet distributions in new physics scenarios, significant computing power is required to deduce the surviving regions of the parameter space of constrained minimal super-symmetric standard model (CMSSM) using Bayesian posterior probability and likelihood function ratio tests.
@@ 5.1 Jet parameters scan
To mitigate computational demands, the study [85] utilizes an MLP as a regressor to learn the mapping from CMSSM model parameters $\theta$ to weak-scale supersymmetric particle masses $m$. The output of the SoftSusy physical package serves as the target output value of the neural network. Approximately 4000 sample points in the parameter space form the training set. With a given set of CMSSM parameters, this MLP model rapidly predicts the corresponding supersymmetric particle mass spectrum, which can then be used to forecast observable distributions at the LHC, including Jet multiplicities and kinematic features. This approach significantly accelerates the process compared to traditional methods. To identify the parameters of a new physics model, [86] trained an MLP using 84 physical observables from the 14 TeV LHC as inputs, many of which involve Jets and their kinematic properties, with the parameters of a supersymmetric model as the desired outputs. The study revealed that with a collider luminosity of 10 $fb^{-1}$, the CMSSM model’s parameters $M_{0}$ and $M_{1/2}$ could be reliably determined with just a 1% margin of error. With a collider luminosity of 500 $fb^{-1}$, additional model parameters such as $\tan\beta$ and $A_{0}$ could also be accurately estimated. In contrast, the conventional approach of minimizing $\chi^{2}$ yielded comparatively inferior results.
Generating collider event samples at the LHC through Monte Carlo simulation can be time-intensive, especially when analyzing detailed Jet structures. While a rapid detector simulation requires only a few minutes, a comprehensive simulation using the GEANT4 framework, as employed by ATLAS and CMS, may necessitate several days. To address this, [87] applied parallel full detector simulations using four parameters—common scalar mass ($m_{0}$), universal gaugino mass ($m_{1/2}$), the trilinear coupling ($A_{0}$), and the ratio of vacuum expectation values ($\tan\beta$)—to produce events including Jets and other final-state objects more efficiently. Two ML models, the MLP and SVM, were employed to learn the correlation between the number of signal events and the CMSSM parameters. The results showed that predicting the likelihood function, which strongly depends on Jet signatures and other observables, could achieve several percent accuracy with just 2000 training samples. Moving on, the paper [88] proposes a MLS (MLS) framework for efficient exploration of multi-parameter supersymmetric models, surpassing traditional methods like MCMC and MultiNest. Utilizing deep neural networks, the MLS incrementally learns parameter space, reducing computational costs while improving target discovery. It integrates HEP packages for precise calculations, including tools like GAMBIT and micrOMEGAs, demonstrating efficiency on toy and CMSSM datasets. Achieving up to 80% sampling efficiency in constrained parameter spaces, MLS outperforms MultiNest under 68% and 95% confidence levels, offering scalability and adaptability for physics model analysis.
@@ 5.2 Jet classification and tagging
Despite treating Jets as images or PC in the calorimeter and exploiting the benefits of DNNs in classification for improved Jet substructure detection, these approaches encounter hurdles. Challenges such as Jet image sparsity and potential precision loss arise from constructing Jet images through pixelation or creating advanced Jet features. In this study [45], a sequential method is employed, utilizing an ordered sequence of Jet constituents as inputs for training. Unlike many prior methods, this approach avoids information loss during pixelization or high-level feature computation. The Jet classification technique achieves a considerable background rejection efficiency operating point for reconstructed Jets with transverse momentum ranging from 600 to 2500 GeV. Moreover, it remains unaffected by multiple proton-proton interactions at levels anticipated during Run 2 of the LHC.
[…]
@@
Multi-Jet classification is a a key task in particle physics aimed at distinguishing between events with varying numbers of Jets. Using ML techniques, such as DNNs, researchers develop classification models to accurately identify these events. Achieving high classification accuracy is crucial for understanding fundamental particle interactions and discovering new physics phenomena in experiments like those conducted at the LHC. The work in [17] present an application of scalable DL to analyze simulation data from proton-proton collisions at 13 TeV in the LHC. The researchers developed a CNN model which utilizes detector responses as two-dimensional images reflecting the geometry of the CMS detector. The model discriminates between signal events of R-parity violating super-symmetry and background events with multiple Jets resulting from inelastic QCD scattering (QCD multi-Jets). With the CNN model, they achieved 1.85 times higher efficiency and 1.2 times higher expected significance compared to the traditional cut-based method. They demonstrated the scalability of the model at a large scale using high-performance computing (HPC) resources with up to 1024 nodes. The authors in [56] proposing an interpretable network for multi-Jet classification using the Jet spectrum, termed S2(R), derived from a Taylor series of an arbitrary Jet MLP classifier function. The network’s intermediate feature is an infrared and collinear safe variables, named C-correlator, estimating the importance of S2(R) deposits at angular scales. It offers comparable performance to CNNs with simpler architecture and fewer inputs. The paper [92] proposes a jet origin identification method for the electron-positron Higgs factory, classifying Jets into 11 categories: 5 quark species, 5 anti-quarks, and gluons. It achieves jet tagging efficiencies ranging from 67% to 92% and charge flip rates between 7% and 24%, utilizing the ParticleNet model. The method benefits jet physics and HEP by enhancing rare Higgs decay measurements. It reduces QCD backgrounds and improves flavor tagging, crucial for Higgs boson property studies. The dataset consists of simulated $\nu\bar{\nu}H,H\rightarrow jj$ events at 240 GeV, generated with a Geant4-based detector simulation. The best reported performance includes a 92% efficiency for $b$-jets and a 7% charge flip rate for charm quarks.